Energy required to accelerate record from 10 metre per second 220 metre per second compared with required to accelerated from zero to 10 M per second inverse is

Grade:10

1 Answers

Aman gohel
6 years ago
Case 1:The amount of energy required is equal to increase in its kinetic energy.Therefore,E1=1/2(m*(220^2-10^2))=>E1=1/2(m*48300)CASE 2:Same rule as above,Therefore,=>E2=1/2*(m*(10^2-0^2))=>E2=1/2*(m*100)Therefore,E1:E2=48300/100=>483:1(Ans)